### v2

# Understanding the chemistry of temperate exoplanet atmospheres through experimental and numerical simulations

11institutetext: 1 LATMOS/IPSL, UVSQ Université Paris-Saclay, Sorbonne Université, CNRS, Guyancourt, France
2 ETH University, Center for Origin and Prevalence of Life, Department of Earth and Planetary Sciences, 8092 Zurich, Switzerland
3 ENS-Paris-Saclay, Gif sur Yvette, France
Understanding the chemistry of temperate exoplanet atmospheres through experimental and numerical simulations
O. Sohier    A. Y. Jaziri    L. Vettier    A. Chatain    T. Drant    N. Carrasco
(Received 03 October 2025; accepted 17 December 2025)
Context: Characterizing temperate exoplanet atmospheres remains challenging due to their small size and low temperatures. Recent JWST observations provide valuable data, but their interpretation has led to diverging conclusions, highlighting the limitations of observations alone. Complementary approaches combining laboratory experiments and photochemical modeling are essential for constraining atmospheric chemistry and interpreting observations.
Aims: This study investigates out-of-equilibrium chemistry in the upper atmospheres of H2-dominated temperate sub-Neptunes enriched in carbon-bearing species (CH4, CO, or CO2). We aim to identify chemical pathways governing the formation and evolution of neutral species and to assess their sensitivity to key parameters such as C/O ratio and metallicity.
Methods: Our approach combines experimental and numerical simulations on H2-rich gas mixtures representative of sub-Neptune atmospheres and spanning a wide range of CH4, CO, and CO2 mixing ratios. We used a cold plasma reactor to simulate out-of-equilibrium upper-atmospheric chemistry. Chemical evolution was tracked by mass spectrometry and infrared spectroscopy. A 0D photochemical model was used to reproduce reactor conditions, guiding interpretation of the key pathways and abundance trends.
Results: We observed the formation of both reduced and oxidized organic compounds. In CH4-rich mixtures, hydrocarbons formed efficiently through methane chemistry, correlating with CH4 concentration and agreeing with models. In more oxidizing environments, particularly CO2-rich mixtures, hydrocarbon formation was inhibited by complex reaction networks and oxidative losses. We find that oxygen incorporation enhances chemical diversity and promotes the formation of oxidized organic compounds of prebiotic interest (H2CO, CH3OH, CH3CHO), especially in atmospheres containing both CH4 and CO2. Atmospheres containing CH4 and CO — which balance carbon and oxygen supply without excessive oxidative destruction — favor efficient production of hydrocarbons and oxidized compounds.
Conclusions: Out-of-equilibrium chemistry plays a key role in the diversification and organic complexification of temperate exoplanet atmospheres. Combining laboratory experiments with photochemical modeling elucidates pathways to hydrocarbon and oxidized organic formation. Studying detectability of these photoproducts with JWST and new high-resolution ground-based instruments is an important focus for future studies.
Key Words.:
exoplanets - atmospheres – chemistry - experimental simulations - numerical simulations
[…]
1 Introduction
Characterizing the atmospheres of temperate exoplanets presents significant challenges due to their small size and relatively cold temperatures, making them difficult to detect and study. These exoplanets, which have an equilibrium temperature below $\sim$500 K, can be classified into three categories: terrestrial planets, super-Earths, and sub-Neptunes. To date, no terrestrial exoplanets, including those in the TRAPPIST-1 system, have been clearly and robustly detected as having an atmosphere (Gillon et al. (2025) for TRAPPIST-1b and c, Piaulet-Ghorayeb et al. (2025) for TRAPPIST-1d, Espinoza et al. (2025); Glidden et al. (2025) for TRAPPIST-1e). On the other hand, sub-Neptunes and super-Earths — which to this day are the most commonly detected types and therefore dominate the population of known exoplanets (Fressin et al., 2013; Luger and Barnes, 2015; Owen and Mohanty, 2016; Van Eylen et al., 2018; Hardegree-Ullman et al., 2019) — have emerged as ideal targets for atmospheric characterization. In particular, those orbiting M dwarf stars have a unique advantage known as the “M dwarf opportunity,” because the smaller size and lower brightness of the host star improve the signal-to-noise ratio during observations. This results in larger transit depths, making it easier to detect and study planetary atmospheres (Cabot et al., 2024; Doyon, 2024). These planets are also particularly intriguing because they have no direct analogs in our Solar System. This raises fundamental questions about the processes that shape their atmospheres, interiors, formation, and habitability (Cabot et al., 2024). In addition, a significant number of planets orbiting M dwarfs reside in their star’s temperate zone (Dressing and Charbonneau, 2015), making the question of their habitability all the more interesting. This habitability is challenged by the unique environmental conditions around M dwarfs, including intense stellar flares, high-energy particle events, and limited ultraviolet flux to drive prebiotic chemistry (Scalo et al., 2007; Ricker et al., 2015; Shields et al., 2016; Rimmer et al., 2018).
Recent advancements in observational techniques, particularly with instruments on the James Webb Space Telescope (JWST), have made these planets favorable targets. In particular colder exoplanets, with temperatures between 500 and 1000 K, tend to have clearer atmospheres, as they are less obscured by clouds and hazes (Morley et al., 2015; Brande et al., 2024; Yang and Hu, 2024; Gao et al., 2021). The advancements have already led to the successful characterization of two temperate sub-Neptune exoplanetary atmospheres, marking significant progress in this field. The analysis of transit spectroscopy data from the atmospheres of K2-18 b (Madhusudhan et al., 2023) and TOI-270 d (Benneke et al., 2024; Holmberg and Madhusudhan, 2024) has highlighted atmospheric compositions dominated by hydrogen and containing a few percent of carbon-bearing species, such as methane or carbon dioxide. Despite these revolutionary new data, their atmospheric composition is still poorly constrained and highly debated.
In the case of K2-18 b, initial JWST observations in the 0.9–5.2 µm range combined with data from NIRISS SOSS and NIRSpec G235H instruments acquired in 2023 suggest the presence of CH4 and CO2 at comparable levels, around 1% each (Madhusudhan et al., 2023). Subsequent studies incorporating non-equilibrium chemistry have led to a revised interpretation, indicating a higher level of CH4 and a lower abundance of CO2 (Wogan et al., 2024). This reinterpretation is supported by a new analysis of the JWST data, which has called into question the initial detection of CO2 as a whole (Schmidt et al., 2025). In 2024 and early 2025, four new transit observations were acquired with the NIRSpec G395H instrument and presented by Hu et al. (2025). By stacking all NIRSpec spectral data obtained to date, they report a significant detection of both CH4 and CO2, with upper limits of approximately 10% and 1% respectively, supporting the hypothesis of a predominance of methane. Meanwhile, initial data were also acquired with the JWST MIRI LRS instrument in the 6–12 µm range, and primary analysis of the data did not provide strong evidence for CH4 or CO2 (Madhusudhan et al., 2025), which do not show clear bands in this range.
The detection of signatures of minor atmospheric constituents also remains largely unconstrained by these low-resolution observations. Non-equilibrium chemistry models help characterize the complete atmospheric composition, but they require observational constraints such as metallicities and C/O ratios, which are also poorly known. They predict the abundance of various components, even minor ones for which signatures remain undetectable in observations. According to these simulations, CH4, CO2, CO, N2, NH3, H2O, C2H4, and C2H6 could be major species, up to the percent range (Jaziri et al., 2025; Luque et al., 2025). Jaziri et al. (2025) also predict that oxidized species such as H2CO, CH3OH, and CH3CHO can reach mixing ratios close to the parts-per-million level. These molecules are of significant prebiotic interest. H2CO is important as a possible precursor to sugars through the formose reaction, which is key for the origin of the RNA or pre-RNA world (Butlerow, 1861; Cleaves, 2008). CH3OH is a key synthon in the RNA world (Mathew et al., 2022). Aldehydes, such as H2CO and CH3CHO, are key precursors of several amino acids through the Strecker synthesis (Strecker, 1850). The atmospheric photochemical production of these oxidized compounds has already been identified as one of the main sources of these compounds in the atmosphere of primitive Earth (Pinto et al., 1980; Wen et al., 1989). It is therefore worthwhile to study their formation pathways in temperate exoplanetary atmospheres.
If we are to overcome the major scientific challenge of characterizing the chemistry of temperate exoplanet atmospheres, a synergy between the observational, modeling, and experimental communities is required. Initiatives of this type are already underway in the community. One example is the attempt to obtain experimental data on the temperature-dependent photo-absorption cross sections of key molecules such as CO2, which provide essential data for atmospheric models of hot exoplanets (Venot et al., 2013, 2018). For the cooler exoplanets, studies have been carried out on the detectability of ions in the thermosphere of sub-Neptunes by combining laboratory experiments, chemical modeling, and simulated observations as well as in the atmospheres of super-Earths by combining laboratory measurements and a 0D photochemical model (Bourgalais et al., 2020, 2021).
A few experiments have been developed for the characterization of the atmospheres of exoplanets, and they mainly focus on the formation and properties of haze and aerosols in the atmospheres of super-Earths and sub-Neptunes (Gavilan et al., 2018; He et al., 2018a, b; Hörst et al., 2018; Moran et al., 2020; Jiménez-Escobar et al., 2025) or hot Jupiters (Fleury et al., 2019, 2020, 2023). Studies focusing on the neutral gas phase chemistry in the atmospheres of cool exoplanets are quite rare. He et al. (2019) conducted laboratory experiments on a wide range of exoplanet atmospheres, with a wide variety of metallicities (100×, 1000×, and 10000× solar metallicity) and temperatures (at 300, 400, and 600 K) and for warmer H2-rich atmospheres (800 K, 100x, and 1000x solar metallicities (He et al., 2020)). They explored the effect of different types of irradiation, using a cold plasma or a UV photon lamp as energy sources. They found that a complex non-equilibrium chemistry occurred, leading to the formation of potential haze-forming precursors. Wang et al. (2025) explored CO2-rich atmospheres at 300 and 500 K, with a 2000x solar metallicity. Under plasma irradiation, they observed the formation of various hydrocarbons and oxygen- and nitrogen-containing species, and they identified reactive gas precursors for haze formation such as C2H4, H2CO, and HCN. However, the analytical diagnosis of all these studies was based solely on residual gas analyzer mass spectrometry (RGA MS), which led to an uncertain identification of some compounds. Reed et al. (2024) studied the UV irradiation of H2S/CH4/N2 gas mixtures, with or without CO2, at an ambient temperature (around 300 K). They analyzed the gas-phase products using gas chromatography equipped with sulfur chemiluminescence detection (GC-SCD), a diagnostic tool less ambiguous than RGA MS. They were thus able to demonstrate the abiotic production of several organosulfur gases, limiting their robustness as biosignatures.
The aim of this work is to propose chemical mechanisms occurring in a wide range of possible atmospheric compositions of temperate sub-Neptunes. In this context, we carried out parallel experimental and 0D numerical simulations to understand the complex out-of-equilibrium chemistry occurring in the upper layers of the H2-dominated atmospheres of temperate exoplanets. These low-pressure atmospheric layers correspond to those probed by JWST observations and were exposed to energetic radiation that triggers photochemical processes. As the atmospheric composition of these exoplanets is unconstrained, we carried out a study on various H2/CH4/CO/CO2 gas mixtures. The C/O ratio (from 0.5 to infinity) and metallicity (from 5 to 50 times the solar metallicity) of these mixtures are variable. The C/O ratio measures the relative abundance of carbon compared to oxygen, while metallicity measures the abundance of elements heavier than hydrogen and helium. To vary these parameters in our H2-dominated gas mixtures, we used CH4, CO, or CO2 as the main component as well as different combinations of these carbon compounds. We chose to focus on these reactants because they allow us to obtain a wide range of possible atmospheres based on the observational data currently available for K2-18 b. Of course, the reactants are relatively simplistic atmospheric analogs that do not take into account other elements such as nitrogen or sulfur, which may also be present in this type of atmosphere. Other oxygen-bearing molecules, such as water, could also be present in significant quantities in this atmosphere, although the upper layers may instead be dried out by cold trap processes (Charnay et al., 2021). Nevertheless, these atmospheric analogs already allow us to cover a wide range of atmospheric possibilities and to identify a great molecular diversity. A key aspect of this work lies in the combination of experiments and numerical simulations, which enables us to extract abundance trends and chemical formation pathways. Identifying key mechanisms enables us to understand the role of different carbon-bearing reactants (CH4, CO, CO2) in organic growth processes (i.e., carbon chains elongation) and to study the effect of oxygen incorporation on chemical diversification.
2 Methods
We studied the chemical evolution of atmospheric analog mixtures when irradiated in a cold plasma setup, mimicking photochemistry in the upper atmosphere. Complementary diagnoses were performed using mass spectrometry (MS) and infrared spectroscopy. In parallel, a 0D photochemical model predicted abundance values and reaction pathways, which we systematically compared with experimental results.
2.1 Mixtures of atmospheric analogs
Different gas mixtures were chosen to study a wide range of potential atmospheres, based on information gathered from observations of K2-18 b. These mixtures are dominated by hydrogen, with carbon-bearing compounds ranging from highly reduced carbon (CH4) to an intermediate state (CO) to fully oxidized carbon (CO2). To enable comparisons with planetary atmospheric models, the C/O ratios and metallicities of our mixtures are provided for reference. It should be noted that our compositions are set by the continuous gas flows in an open system. The gas residence time, i.e., the time spent by the gas in the irradiated medium, is set by the pressure and gas flow in our experiments and thus differ from the residence time found in atmospheric models. Thus, the C/O ratios given are enriched in CH4, CO and CO2, compared to atmospheric models. All mixtures are summarized in Table 1.
The most reduced atmospheres are composed of H2 and CH4. The relative proportions of H2 and CH4 vary in the experiments (see Table 1 for the relative mixing ratios). We refer to these mixtures as reduced mixtures. This type of atmosphere helps us characterize organic growth in the most reducing environment.
The more oxidized atmospheres are composed of H2 and CO or CO2. The relative mixing ratios are shown in Table 1. We refer to these mixtures as oxidized mixtures. This type of atmosphere allows us to explore the balance between adding carbon and oxygen atoms to the system.
The composite atmospheres combine the conditions of the reduced and oxidized atmospheres described above. They are composed of H2, CH4, CO and/or CO2. A wide range of oxidation states is studied, as shown in Table 1.
Table 1: Experimental compositions in percentages.
Reduced mixtures Oxidized mixtures CH4/CO/CO2 mixtures
H2 99 97 95 90 99 97 95 90 99 97 95 90 98 94.5 94 94 90 98 94.5 94 94 90 97,5 89,5
CH4 1 3 5 10 1 5 1 5 5 1 5 1 5 5 1 5
CO 1 3 5 10 1 0.5 5 1 5 1 5
CO2 1 3 5 10 1 0.5 5 1 5 0.5 0.5
C/O ratio $\infty$ $\infty$ $\infty$ $\infty$ 1 1 1 1 0.5 0.5 0.5 0.5 2 11 1.2 6 2 1 5.5 0.6 3 1 1.25 1.75
Metallicity (compared to solar) 5.6 14.8 22.1 35.1 9.2 22.6 31.7 45.4 13.6 30.2 40.0 53.0 13.7 24.8 33.8 27.2 41.0 17.5 26.2 41.5 29.7 46.6 18.9 42.9
2.2 Atmospheric simulation chamber
Experimental simulations of the upper atmospheric layers of temperate exoplanets were carried out using an experimental cold plasma device, the PAMPRE reactor (for Production d’Aérosols en Microgravité par Plasma REactif (Szopa et al., 2006)). This setup was initially developed to simulate the atmospheric chemistry of Titan, and the production of aerosol analogs (tholins) created by photochemistry in this atmosphere (Szopa et al., 2006; Quirico et al., 2008; Sciamma-O’Brien et al., 2010; Carrasco et al., 2012; Mahjoub et al., 2012; Fleury et al., 2014; Dubois et al., 2019, 2020). The use of this experimental platform has been extended over the years to include other atmospheres, such as early Earth (Fleury, 2015; Fleury et al., 2017) or Pluto (Jovanovic, 2021). More generally, the versatility of the experimental setup means we can study any temperate atmospheric chemical composition, which make it suitable for studying exoplanets. Initial research has already been carried out on organic aerosols produced in the atmospheres of Earth-like exoplanets (Gavilan et al., 2017, 2018).
PAMPRE induces a cold capacitively coupled radiofrequency plasma in a gaseous mixture, outside of thermodynamic equilibrium. This reproduces the energetic processes occurring in the upper layers of a temperate atmosphere exposed to energetic stellar particles and UV irradiation. It has been previously shown that the energy distribution of electrons in a nitrogen-dominated plasma reproduces the energy distribution of solar photons fairly accurately, with a slightly higher contribution in the VUV (Szopa et al., 2006). The stellar spectra of M dwarfs — around which the temperate exoplanets that interest us orbit — have higher UV radiation contributions than the Sun (France et al., 2013). This experimental setup is therefore appropriate for simulating the energetic environment of these exoplanetary atmospheres. Although our plasma is dominated by hydrogen, which may slightly alter the energy distribution of electrons, it remains adequate. These energetic processes trigger an imbalance in chemistry, dissociating gas molecules and allowing fragments to reorganize into new species. In specific cases, products even aggregate into solid particles, the tholins.
The setup consists of a cylindrical stainless steel reactor, 40 cm high and 30 cm in diameter, connected to a quadrupole mass spectrometer (QMS). High-pressure gas cylinders are connected to the setup via an injection manifold, enabling gas mixtures to be introduced into the main chamber. MKS mass flow meters, rated at 5, 10 and 100 standard cubic centimeters per minute (sccm) with a full-scale accuracy of 1%, control the injection rates. The plasma is confined in a stainless steel cage measuring 13.8 cm in diameter and 5 cm in height, pierced with small holes to allow the gas to pass through. The top of this cage is the cathode, measuring 12.6 cm in diameter, polarized and powered by a 13.56 MHz generator. The gas mixture is injected into the reactor from the top of the chamber through the cathode, which is equipped with three superimposed grids to ensure a continuous, homogeneous injection flow. Gas flow rate is set at 40 sccm. The central grid, with a mesh size in the order of the Debye length ($\sim$0.1 mm), prevents the plasma from flowing back into the gas injection system. The gas mixture is extracted from the bottom of the chamber by a primary vane pump (Adixen by Pfeiffer Vacuum Pascal 2015 SD). This enables an open continuous-flow system, at 0.98 mbar and at ambient temperature ($\sim$300 K). A window allows visual monitoring of the plasma. Plasma discharge is only used in continuous mode in these experiments, with a power of 30 W. Between each experiment, the reactor is pumped with a turbomolecular pump (Adixen ATP 80) to achieve a secondary vacuum of around 1.5×10-6 mbar. A schematic diagram of the experimental setup is shown in Figure 1.
Figure 1: Scheme of the plasma reactor PAMPRE.
2.3 Mass spectrometry
Mass spectrometry is used as an analytical method to determine the molecular species produced. PAMPRE is equipped with a Hiden EQP system, which is used in residual gas analysis (RGA) mode, since only neutral species are considered in this study. A collector head is used to sample gaseous species through a 100 µm diameter orifice. The pressure inside the mass spectrometer is kept below 5x10-6 mbar. To study neutral species, an internal ionizer is used, with electron energy set at 70 eV. The resulting ions are then filtered and focused before entering the triple-filter mass analyzer, which detects masses from 1 to 200 m/z (mass-to-charge ratio) with a resolution of 1 atomic mass unit (amu). Finally, the ions arrive at an ion-counting detector — a secondary electron multiplier.
Several molecular fragments at different m/z values are generated and form a ”fingerprint” specific to each molecule and incident electron energy, known as the fragmentation pattern. A full mass spectrum thus corresponds to the superposition of the fragmentation patterns of all the species present in the analyzed gas mixture. Reference fragmentation patterns were measured directly in the lab for the reactive compounds, i.e., H2, CH4, CO and CO2. The reference fragmentation patterns of the products are extracted from the NIST Chemistry WebBook [^1] https://webbook.nist.gov/chemistry/ (Linstrom and Mallard, 2001).
All m/z over a selected mass range are scanned in BAR mode (Batched Analysis of Repetitions). This produces a histogram of peak intensities. Each mass spectrum contains 20 accumulated scans, over a mass range from 1 to 100 amu. A mass spectrum is taken at the initial state (spectrum OFF), when the gaseous mixture is injected inside the reactive chamber at 0.98 mbar, while the plasma discharge is still switched off. The plasma is switched on and after five minutes, a spectrum is taken (spectrum ON). To ensure that we had reached a stable steady state, we performed multiple ion detection when we activated the plasma. This allowed us to track several m/z ratios over time, thereby monitoring consumption and production and determining when we reach a stable state. In all our experiments, we observed stabilization after a maximum of 200 seconds.
The evolution of the gas composition can be seen by our plotting of the superposition of the two ON and OFF spectra (Figure 2). Our plotting of the difference of intensity between these two spectra (ON-OFF) reveals the production and destruction of the chemical species.
Figure 2: Mass spectrum of a gas mixture of 95% H2 plus 5% CH4 before (gold) and during irradiation (blue).
The measured mass spectra are processed first by removing the noise from the data: all intensities below 10 counts/s are removed, as the instrument’s sensitivity is 106 counts/s maximum. Each spectrum is then normalized by its maximum intensity. In the direct-flow experiments, the maximum intensity is always at m/z 2 amu, corresponding to H2. In trap-and-release experiments, the maximum intensity depends on the composition. Mean values and standard deviation at 95% are calculated from the 20 cycles for each m/z. If the difference between the average intensity and the standard deviation is below 10-6, it is considered as noise and removed from the data.
While the reconstruction of a mass spectrum with fragmentation patterns works well with a few compounds, it becomes much more difficult to analyze and identify species in more complex mixtures with dozens of products. This is due to the superposition of several fragmentation patterns and the variability associated with these patterns — which depends on the geometry of the ion source, the mass-dependent filtering, and detection efficiency of the instrument. We therefore need a complementary analytical method in order to make unambiguous identifications.
2.4 Infrared spectroscopy
Infrared spectroscopy was used to help identify species in CH4/CO/CO2 mixtures. The PAMPRE setup was connected to a multipass infrared white cell with an optical beam length of 10 m. The cell was positioned in the sample compartment of an FTIR spectrometer (Nicolet 6700, Thermo Fisher).
Gas mixtures flow directly through this cell to characterize the dissociation rate of reactants and the formation of major products. However, the concentration of minor products is too low relative to the instrument’s thresholds to allow their identification. To overcome this limit, we use a cold trapping strategy to concentrate our products before IR identification. A glass trap installed at the outlet of the system is immersed in liquid nitrogen (78 K) (see Figure 1), ensuring the entire flow passes through it, condensing most of the gases. H2, CH4 and CO, with condensation temperatures below 78 K, are not trapped, which helps to concentrate minor products. However, precise quantification is not possible due to this bias-inducing differentiated condensation, but robust product identification is already an important step forward. The irradiated gas stream is allowed to pass through the trap for 20 min; trapping is then stopped, the trap is heated to sublimate the condensed species, the gases are released into the multipass cell, and heated to 70°C so that water vapor does not condense on the mirrors. The pressure is measured inside the cell and it does not exceed a few millibars. Since there is no longer any irradiation, we remain at low pressure and do not reach very high temperatures, so we expect the kinetics to be slow and no significant new reactions to occur. Infrared spectra are measured through the software OMNIC (Thermo Scientific), with a resolution of 0.5 cm-1using the Graymann-Harris apodization function Harris (1978). 150 scans were accumulated for each measurement, averaged, and combined into a spectrum.
Preprocessing the spectra, using OMNIC software, transforms the single-beam data acquired into transmittance or absorbance spectra, by dividing them by blank spectra previously acquired under high vacuum conditions (less than 10-5 mbar) and by applying a manual baseline correction. The IR spectra were compared with those calculated at the experimental pressure (1 mbar) and temperature (340 K), based on the opacities of several species (Figure 12). The opacities are taken from the databases DACE  [^2] https://dace.unige.ch/opacityDatabase/? and HITRAN (Gordon et al., 2017). However, these identifications were limited by the availability of molecules — organic compounds with three or more carbon atoms are not included. For more complex molecules, identification is performed using the NIST database, but this does not allow for temperature and pressure dependence to be taken into account.
2.5 Atmospheric numerical simulations
To model the chemistry occurring in the reactor chamber, we used the 0D photochemical model ReactorUI (Peng et al., 2014; Pernot, 2023), as described in Peng et al. (2014). This complementary approach to experimental characterization helps to identify the species produced and to elucidate the main chemical pathways leading to their formation (Bourgalais et al., 2020).
The model reproduces experimental conditions, taking into account chamber geometry, gas flows and the distribution of incoming plasma energy, which has been approximated using solar spectra. This is a significant assumption that can influence abundance but still allows pathways to be identified. After adjusting all the parameters, the integration system converges to a steady state.
The chemical network was originally developed for the CHN atmosphere of Titan with neutral-neutral reactions (Hébrard et al., 2007, 2009; Dobrijevic et al., 2016), ions (Carrasco et al., 2008; Plessis et al., 2012) and photo-processes (from the Leiden (Heays et al., 2017) and SWRI (Huebner and Mukherjee, 2015) databases). For the studies of the formation of more complex oxygenated molecules, such as H2CO and CH3OH, we extended the chemical network with that of CHO VULCAN (Tsai et al., 2021), specific to sub-Neptunes, based on Venot et al. (2020).
Simulations of each experimental composition were run. Considering chemical rate uncertainties with a Monte Carlo approach, we perform 500 runs for each configuration. This Monte Carlo approach to characterizing uncertainties has already been used in several atmospheric models of Solar System objects, such as Triton (250 runs, Benne et al. (2022)), Titan (500 runs, Hébrard et al. (2009)) or Neptune (1000 runs, Dobrijevic et al. (2010)). As a result, we evaluate the final species abundance values with its uncertainty. These simulations also provide information on the contributions of specific reactions to the formation or consumption of a given molecule, enabling reaction paths to be extracted.
3 Results
3.1 Identification of main compounds
Table 2 details the value of the main IR and MS features of the major products. Figure 12 shows the reference IR spectra extracted from the HITRAN database for the molecules containing up to two carbon atoms considered in this study.
3.1.1 Reduced organic compounds
We observe the formation of reduced organic compounds, composed only of carbon and hydrogen atoms, forming chains with up to six carbon atoms. In MS, groups of peaks form at m/z up to 80 when a 95% H2 + 5% CH4 gas mixture is irradiated (Figure 2). The specific fragmentation patterns (FP) of reduced compounds with two and three carbon atoms (C2 and C3) are shown in Figure 11. These patterns are normalized to the mass spectrum measured in the irradiated gas mixture of 95% H2 + 5% CH4, and directly compared with the corresponding mass spectrum. In the case of reduced atmospheres, C2 identification using MS alone is possible, particularly for C2H2 and C2H6, which exhibit strong characteristic peaks. Infrared spectroscopy is used to determine the nature of the chemical bonds between carbon atoms, distinguishing between single, double and triple bonds. The IR features used to identify acetylene (C2H2), ethene (C2H4), propyne (C3H4) and propene (C3H6) are shown in Figure 3.
3.1.2 Oxidized organic compounds
Using IR, we could see in CO and CO2-rich mixtures the robust formation of the oxidized organic compounds methanol (CH3OH), formaldehyde (H2CO) and acetaldehyde (CH3CHO). The features used for their identification are shown in Figure 4, at 1033 cm-1(9.68 µm) for CH3OH and at 1745 cm-1(5.73 µm) for the aldehydes (H2CO and CH3CHO). In MS, the production of CH3OH is identified by the appearance of the characteristic peaks at m/z 31 and 32. The fragmentation patterns of H2CO and CH3CHO overlap with those of C2H6 and C3H6, which makes their identification only through MS a source of uncertainty.
3.1.3 Other oxidized compounds
Depending on the degree of oxidation of the reactants, the formation of other oxidized molecules is identified. H2O is identified in MS with peaks at m/z 17 and 18, and presents wide features in IR. CO is identified with a specific peak at m/z 28 in MS, but is not trapped for IR measurements as discussed in section 2.4. CO2 is identified in MS with a peak at m/z 44, and in IR with a narrow peak at 668 cm-1(14.97 µm) and a broad one from 2200 to 2400 cm-1(4.17-4.55 µm).
(a) Infrared feature at 730 cm-1(13.70 µm), attributed to C2H2.
(b) Infrared feature at 950 cm-1(10.53 µm), attributed to C2H4.
(c) Infrared feature at 2150 cm-1(4.65 µm), attributed to C3H4.
(d) Infrared feature at 912 cm-1(10.96 µm), attributed to C3H6.
Figure 3: Infrared features of the reduced organics.
(a) Infrared feature at 1033 cm-1(9.68 µm), attributed to CH3OH.
(b) Infrared feature at 1745 cm-1(5.73 µm), attributed to aldehydes.
Figure 4: Infrared features of the oxidized organics.
3.2 Abundance and formation trends
3.2.1 Reduced organic compounds
(a) H2 + CH4
(b) H2 + CO
(c) H2 + CH4 + CO
(d) H2 + CO2
(e) H2 + CH4 + CO2
Figure 5: Mass spectra acquired for different gas mixtures while plasma is ON.
(a) Reduced versus oxidized mixtures
(b) CO-rich versus CO2-rich mixtures
(c) CH4/CO versus CH4/CO2 mixtures
(d) CH4/CO versus CH4/CO/CO2 mixtures
Figure 6: Predicted abundance values with 0D simulations of the two-carbon atoms reduced organics in the different gas mixtures.
The formation of reduced organics is observed under all experimental conditions. We noted the appearance of carbon chain features both in MS and IR as presented in section 3.1, shown in Figures 2 and 3, and detailed in Table 2.
The organic growth is most significant in reduced mixtures, and it increases with methane concentration and hence with metallicity, as expected and already highlighted by previous experimental studies (Hörst et al., 2018). The mass spectra show an increase in the intensity of all peaks with increasing methane content, as shown in Figure 5(a). We observed peak groups characteristic of compounds comprising up to four carbon atoms (C4), at m/z 50-60. Although it is difficult to quantify the species precisely, the relative intensities of the peaks relating to C2 compounds appear to be of comparable magnitudes, reflecting similar concentrations. Based on the 0D simulations performed with ReactorUI, we observe that the abundance of C2H2 is approximately two orders of magnitude lower than that of C2H4 and C2H6, and decreases with increasing CH4 concentration (Figure 6(a)). In contrast, the abundance of C2H4 and C2H6 increase with increasing methane concentration.
The formation of reduced organics also occurs in oxidized mixtures, although it is more limited than in reduced mixtures, as previously highlighted in particular for CO2-rich mixtures (Trainer et al., 2004). It is observed in MS, as we find lower intensities of C2 (m/z 20-30) and C3 (m/z 30-40) specific peak groups in CO and CO2-rich mixtures (Figures 5(b) and 5(d)) than in reduced mixtures (Figure 5(a)). In Figures 5(b) and 5(d), we do not observe the C4 specific peak groups at m/z 50-60 that were present in Figure 5(a), reflecting inhibited organic growth. No selectivity is observed in oxidized mixtures between the C2 compounds, with simulated abundance values similar for C2H2, C2H4 and C2H6 (Figure 6(a),6(b)).
The organic growth is predicted by the 0D-simulations to be more efficient with CO-rich mixtures than with CO2-rich mixtures, as shown in Figure 6(a) where we see that the abundance values of two-carbon chains are higher with CO than with CO2. This is not obvious with the MS results presented in Figures 5(b) and 5(d), where we see similar relative intensities of the characteristic m/z ratios of compounds C2 and C3. However this is supported by the IR data, as we observe the appearance of species absorption peaks of three-carbon chains only in CO-rich mixtures (CO and CH4/CO mixtures). For example, we only see the 912 cm-1(10.96 µm) peak attributed to C3H6 clearly appearing in CH4/CO mixtures, as shown in Figure 3(d) (solid lines).
For CO-rich mixtures, the organic growth increases with CO concentration. This trend is apparent in the mass spectra in Figure 5(b), where we see the intensities of the C2 and C3 peaks increase with CO concentration. This is confirmed by the 0D simulations performed with ReactorUI, as shown in Figure 6(a). The predicted abundance of the C2 compounds is higher at a concentration of 10% of CO than at a concentration of 1%.
For CO2-rich mixtures, this trend is reversed: the organic growth is limited at higher CO2 concentrations. In the mass spectra, the decrease in the m/z 26 peak — characteristic of C2H2, at a CO2 concentration of 10% — reflects this limitation. The 0D simulations support this trend, with the predicted abundance values of C2 compounds being higher at a concentration of 1% CO2 than at a concentration of 10%, as shown in Figure 6(a).
When H2, CH4 and an oxidized compound — CO or CO2 — are merged, the efficiency of organic growth is a hybrid of the two previous cases. It is increased compared to pure oxidized mixtures, and slightly decreased compared to pure reduced mixtures. In the mass spectra, characteristic peak groups of compounds C2 and C3 are seen in Figures 5(c) and 5(e), at higher intensities than in the oxidized mixtures presented above, but lower than in the reduced mixtures. Figure 6(b) shows the effect of adding methane to the atmospheres, particularly for the formation of C2H4 and C2H6: the gain is two to six orders of magnitude, selectively favoring the formation of C2H4 and C2H6 over C2H2. The more oxidized the composite atmosphere is, the less efficient the formation of reduced organics is. This relationship is visible in Figure 6(c), where the predicted abundance values for the 94% H2 + 5% [ CO or CO2 ] + 1% CH4 mixtures are the lowest. We also observe in Figure 6(d) that when 0.5% CO2 is added to CH4/CO mixtures, the reduced organics abundance values decrease.
3.2.2 Oxidized organic compounds
The formation of the oxidized organic compounds methanol, formaldehyde and acetaldehyde is identified through IR features shown in Figure 4, only in the atmospheres with CO or CO2. The trapping method, which concentrates the products so that they can be seen in IR, induces a bias that prevents robust quantification directly from the IR results. However, the 0D model allows us to extract trends for the abundance of these oxidized organic compounds, as a function of gas mixtures, shown in Figure 7.
(a) Oxidized mixtures
(b) CO-rich versus CO2-rich mixtures
(c) CH4/CO versus CH4/CO2 mixtures
(d) CH4/CO versus CH4/CO/CO2 mixtures
Figure 7: Predicted abundance values with 0D simulations of the oxidized organics in the different gas mixtures.
Figure 7(a) shows that the relative abundance values of the three oxidized compounds for CO-rich mixtures are always higher than for CO2-rich mixtures. It also highlights that the growth of these compounds is limited with higher concentrations of CO2, while it is enhanced with high concentrations of CO. The most abundant oxidized organic product observed with 0D simulations is H2CO (Figure 7). It is produced in the same abundance range as reduced organic compounds such as C2H4 and C2H6, and even above the abundance of C2H2 (Figure 6(c)). CH3OH is much less produced — below the ppb range.
When CH4 is added to the atmosphere, the production of oxidized organics is significantly increased (Figure 7(b)), and enhanced at higher concentrations of both CO and CO2. In this case, we observe that the formation of H2CO and CH3OH becomes more efficient with CO2 than with CO. We observe in Figure 4 that for a given concentration (same color), the intensity of the feature always appears to be greater in H2/CH4/CO2 mixtures (dotted lines) than in H2/CH4/CO mixtures (solid lines). This is confirmed by the results of the 0D simulations, where higher abundance values for H2CO and CH3OH are observed in CO2-rich mixtures than in those that are CO-rich (Figure 7(c)). On the other hand, the abundance values of CH3CHO remains higher with CO than with CO2. The combination of CH4, CO and CO2 increases the abundance of all oxidized organics of about one order of magnitude (Figure 7(d)).
3.3 Formation mechanisms
From the 0D simulations performed with ReactorUI, we can determine the key reactions involved in the production and destruction of each chemical compound. This allows us to reconstruct the chemical networks involving the molecules that interest us in our study.
3.3.1 Reduced organic compounds
Figure 8(a) shows the formation pathway of the C2 compounds from the reactant CH4 in reduced mixtures. From the dissociation of CH4, several intermediates are produced, which recombined to form the C2 compounds. C2H6 is mainly formed by the combination reaction of 2 CH3 compounds, C2H4 is mainly produced by the reaction of CH with CH4, and C2H2 mainly comes from the combination of two triplet methylene radicals 3CH2. The radical reaction required for the production of C2H2 — which is approximately three orders of magnitude slower than the combination reactions forming C2H4 and C2H6 — could explain its lower abundance predicted by 0D simulations. The relative contributions of these reactions in the formation of the reduced organics remains more or less the same as the concentration of CH4 increases. That is coherent with the fact that the higher the proportion of CH4 in the atmosphere — and therefore the higher the metallicity — the higher the final concentrations of reduced organic compounds.
This chemical reactions network (Figure 8(a)) is consistent with the well-known methane chemistry and hydrocarbon formation studied in the reduced atmospheres of our Solar System, such as Titan (Nixon, 2024), Jupiter (Moses et al., 2005) or Saturn (Moses et al., 2000), as well as with predictions made for exoplanets with H2 dominant atmospheres (Adams et al., 2022). This validates our approach, and gives us confidence for the study of our more oxidized atmospheric ranges.
Figure 8(b) shows the formation pathway of reduced organics from CO and CO2. Organic growth is inhibited compared to reduced mixtures, as has been previously studied in the literature — particularly in the CO2 case by Trainer et al. (2004). It requires passing through more intermediates than in the reduced atmospheres, which limits the formation of reduced organic compounds. The formation of C2H4 and C2H6 requires the prior formation of the CH3 intermediate, which in turn is produced by two consecutive intermediates (C and 3CH2) after CO dissociation. This long reaction path explains the lower abundance of hydrocarbons compared to the reduced atmosphere. Furthermore, the dissociation of CO, which provides a source of carbon atoms for organic growth, is at the same time a source of exited oxygen radicals, mainly in the O(3P) state, which will oxidize some of the carbonate radical intermediates and reduced organics (Figure 9). For example, 3CH2 reacts with O(3P) to reform CO. However, as the concentration of CO in gas mixtures increases, there is an overall rise in the formation of reduced organic compounds, showing that with CO, the effect of increasing the carbon source dominates over the loss of intermediates by oxidation for the formation of the organics.
For CO2-rich mixtures, the difference in efficiency in the formation of reduced organic compounds compared to CO-rich mixtures can be linked with the fact that it first requires the dissociation of CO2 before producing the intermediates. This produces CO and an exited oxygen radical, mainly in the O(1D) state. The conversion of CO2 into CO is not total, which explains why the source of carbon is less abundant than in the case where CO is a reactant. In addition, the production of oxygen radicals is doubled, as a first radical is produced from the dissociation of CO2 and a second from the dissociation of CO. The oxidation of carbonate intermediates and reduced organics is more important. For example, the reaction of C2H2 with OH radical, which is a minor loss reaction in the case of CO, becomes dominant in the case of CO2. This is related to the fact that the formation of OH radicals is more efficient from CO2 than from CO, thanks to the following combination reaction, which is an order of magnitude faster with CO2 than with CO:
$O^{1}D+H_{2}\longrightarrow OH+H.$ (1)
The main pathways of loss through oxidation are shown in Figure 9. These combined effects explain why the abundance of reduced organic substances in CO2-rich gas mixtures is much lower than in CO-rich mixtures, and why the abundance of reduced organic compounds decreases with increasing CO2.
As CH4 is a reactant in the CH4/CO/CO2 mixtures, the reaction pathway presented in Figure 8(a) also explains the organic growth. When atmospheres are more oxidized, we can observe the same mechanisms that are distinctive of fully oxidized atmospheres, as developed in Figure 8(b), but with a limitation in organic growth due to the oxidation of reaction products and intermediates.
In summary, organic growth efficiency is best in CH4-rich mixtures, as expected, and is significant in oxidized mixtures, although more limited. In mixtures rich in CO2, oxidative loss reactions become predominant and limit hydrocarbon formation, even at high metallicity values. In CO-rich mixtures, the effect of the carbon source partially compensates for the inhibitory effect of the oxygen source, and we observe that organic growth efficiency increases with metallicity, as in reduced gas mixtures.
Figure 8: Formation pathways of reduced organics with main (solid arrows) and minor (dotted arrows) reactions started from a reduced reactant (CH4 8(a)) or an oxidized one (CO or CO2 8(b)).
Figure 9: Main pathways of loss through oxidation of reduced organic compounds (red) superimposed on their formation pathways (light gray) presented in Figure 8(b).
3.3.2 Oxidized organic compounds
The main formation pathways of H2CO and CH3CHO are different depending on the oxidized reactant, as shown in Figure 10(a) and 10(c), while for CH3OH it is similar for CO and CO2. All these reaction paths involve reduced organic intermediates, such as CH3 and 3CH2, for which the formation efficiency in CO or CO2 mixtures is described in section 3.2.1. This explains the formation trends observed, i.e., formation efficiency is better with CO than with CO2, enhanced at high CO concentrations and limited at high CO2 concentrations. The formation of CH3OH and CH3CHO in mixtures of CO and CO2 requires at least one more intermediate than the formation of H2CO. This makes the formation of H2CO up to six orders of magnitude faster, explaining its predominance.
The main loss pathways of oxidized organic compounds could provide an additional explanation for the predominance of H2CO over CH3OH and CH3CHO. Indeed, in all gas mixtures, the loss of H2CO is mainly due to the following reactions:
$H_{2}CO+H\longrightarrow HCO+H$ (2)
$H_{2}CO+h\nu\longrightarrow CO+H_{2}.$ (3)
The reaction of H2CO with H is almost two orders of magnitude faster than its photodissociation in the 0D model. HCO and CO are themselves involved as reactants in the formation of H2CO. In this way, H2CO is self-sustaining, which explains its high abundance. In contrast, CH3OH can be photo-dissociated through
$CH_{3}OH+h\nu\longrightarrow H_{2}CO+H_{2},$ (4)
which forms H2CO again.
The addition of CH4 in the gaseous mixture enhances all the productions, since the formation pathways of the reduced intermediates are much shorter starting from the reactant CH4 than from CO or CO2. With this limiting effect removed, the production of H2CO and CH3OH from CO2 is more efficient than with CO. This is because CO2 is a direct reactant in the formation of H2CO, whereas with CO, the intermediate O(1D) is required (Figure 10(a)), which is a minor product during the dissociation of CO. For the formation of CH3OH, the same reaction pathway is followed from CO or from CO2 (Figure 10(b)), requiring the intermediate O(1D), a major product during the dissociation of CO2. However, since the formation of CH3CHO from CO2 requires one more intermediate than from CO (Figure 10(c)), it remains more efficient with CO than with CO2.
In summary, the efficiency of the formation of oxidized organic compounds is greater in H2/CO mixtures and increases with metallicity. In H2/CO2 mixtures, formation of oxidized organic compounds is limited and decreases with the metallicity. However, the combination of CH4 with CO or CO2 significantly increases the formation of oxidized products, in correlation with metallicity. The efficiency of H2CO and CH3OH formation is therefore higher in CH4/CO2 mixtures than in CH4/CO mixtures.
(a) Formation pathways of H2CO
(b) Formation pathways of CH3OH
(c) Formation pathways of CH3CHO
Figure 10: Formation pathways of oxidized organics. The formation pathways of the intermediates CH3 and 3CH2 from CO or CO2 are detailed in Figure 8 and are simplified here by long dotted gray arrows. The main (solid arrows) and minor (short dotted arrows) reactions are represented.
4 Discussion
4.1 Roles of methane, carbon monoxide, and carbon dioxide in organic growth
Methane plays a predominant role in organic growth, as demonstrated in our experiments, in which we observed a positive correlation between gas-phase hydrocarbon production and methane concentration. This is consistent with previous experiments in the literature, which have highlighted the major role of methane in the growth of solid particles (tholins) when N2/CH4 Titan-type mixtures are irradiated (Hörst and Tolbert, 2013; Hörst et al., 2018; Moran et al., 2022; Jovanović et al., 2020; Jovanovic, 2021). The production of the solid phase in Titan-type mixtures is optimized between 3 and 5% of CH4 (Sciamma-O’Brien et al., 2010). In this study, in H2-dominated mixtures up to 10% of CH4, we observe continuous growth of hydrocarbons in the gas phase, without an obvious optimal concentration. This optimum would possibly occur at higher proportions of CH4.
While methane is clearly identified as a major driver of organic growth, the effect of oxidized reactants is not as clearly correlated. At first glance, we identified an inhibitory effect due to the atomic oxygen source represented by the oxidized reactants, thus limiting the growth of long carbon chains. This is even more pronounced with CO2, which provides twice as much oxygen. This effect has also been identified in the production of solid particles in N2/CH4 mixtures, where the addition of CO alters the kinetics of methane, delaying the appearance of solid particles and reducing their production rate (Fleury et al., 2014; He et al., 2017; Moran et al., 2022).
However, CO and CO2 are also sources of carbon and can contribute to organic growth. Trainer et al. (2006) showed that adding CO2 to a N2/CH4 mixture could increase particle production up to a CH4/CO2 ratio of 1 before decreasing. This effect is particularly evident in our CH4-free mixtures, where hydrocarbon production is still observed, although at lower levels than in the case where CH4 is present. The addition of CO to N2/CH4 mixtures, although reducing the production rate of tholins, increases their size (Fleury et al., 2014; Hörst and Tolbert, 2014). In our study, we observe a positive correlation between the formation of organic compounds in the gas phase and the concentration of CO. That highlights that with CO, the effect of increasing the carbon source dominates over the loss by oxidation. In the case of a low methane concentration (1%), Jovanović et al. (2020) observed a higher production rate in the presence of 500 ppm CO than in its absence, highlighting the important role that CO can play as a carbon source when methane is less dominant.
Furthermore, the addition of oxygen atoms increases the diversification and organic complexity of photochemical products. We observe the appearance of oxidized organic compounds of prebiotic interest in our CO and CO2-rich mixtures (CH3OH, H2CO, CH3CHO). It has also been shown that adding CO to Titan-type mixtures improves the reactivity of the gas phase (Yang et al., 2025; Fleury et al., 2014), leading to the incorporation of oxygen in both the gas and solid phases (He et al., 2017). Similarly, the addition of CO2 leads to more oxidized solid compounds (Trainer et al., 2004). Since CO2 is a more important source of oxygen than CO, one would expect oxygen incorporation to be better. In our study, this corresponds to the observation that oxidized organic compounds are better produced in the gas phase with CO2 than with CO when combined with CH4 and therefore when organic growth is not too limited. For the solid phase, Drant et al. (2025) observed in N2/CH4-rich exoplanetary haze analogs that the characteristic absorption bands of oxygenated functions (O–H, C=O) are clearly visible in analogs produced with CO2, but not or only very weakly in those produced with CO alone.
4.2 Photochemistry of hydrocarbons in temperate methane-rich exoplanet atmosphere
Experiments conducted in reduced gas mixtures reveal similar formation rates for C2 compounds, preventing any definitive conclusions about selectivity. Simulated abundance using the 0D model also predicts comparable values for these molecules, but with a slight selectivity observed for C2H6 and C2H4 relative to C2H2. This contrasts with the hydrocarbon photochemistry observed in the gas giants Jupiter and Saturn, where detected abundance is depleted in C2H4 and enriched in C2H6 and C2H2 (Moses et al., 2000, 2005; Nixon et al., 2010). These planets have hydrogen-dominated atmospheres with significant helium concentrations and a small fraction of methane (around 0.1%). They are cooler — around 150 K at 1 mbar.
The low abundance of C2H4 is explained by its high photodissociation rate at Lyman-$\alpha$, producing C2H2, and by its reaction with H, which recycles it into CH4 (Gladstone et al., 1996). In contrast, C2H6 is shielded by methane and thus remains stable at Lyman-$\alpha$. In our study, the energy distribution of electrons in our plasma covers a much wider range than that of photons in planetary atmospheres, which are largely concentrated around Lyman-$\alpha$. Consequently, the wavelength selectivity of photodissociation is not reproduced, which likely contributes to the differences observed in the relative abundance of C2 compounds between our experiments and planetary atmospheres.
Additionally, C2H2, as a highly unsaturated compound, would normally be unstable in an H2-dominated environment. However, it is efficiently recycled via the recombination reaction
$C_{2}H+H_{2}\longrightarrow C_{2}H_{2}+H.$ (5)
We noticed that this reaction is negligible in our reduced gas mixtures, whereas it becomes significant in oxidized mixtures. Since CO and CO2 are sources of carbon but create a less reducing environment than CH4, this increases the degree of unsaturation of the products. This has also been demonstrated for the production of solid products (He et al., 2017; Yang et al., 2025). This corresponds to the fact that we observe a depletion of C2H2 in mixtures containing only CH4, whereas there is no apparent selectivity in mixtures containing CO and CO2.
Temperature differences may also play a role in C2 relative abundance. Adams et al. (2022) suggests that warmer H2-rich atmospheres, compared to Titan, favor dynamic chemical cycles involving C2H2, C2H3, and C2H4 through reactions with atomic hydrogen. These processes can reduce C2H6 and C2H2 concentrations while increasing the relative abundance of C2H4.
4.3 Detectability of photoproducts
4.3.1 Hydrocarbons
The C2 compounds appear to be relatively abundant and efficiently produced in our experiments, which are dominated by out-of-equilibrium chemistry. Similarly, a 1D non-equilibrium chemical model fitting the JWST observational data of K2-18 b predicts that C2H2, C2H4 and C2H6 could reach volume mixing ratios of up to a few percent in the upper atmosphere (at 1 mbar) (Jaziri et al., 2025). C2H2 and C2H4 are in principle detectable with JWST, using combined observations from the NIRSpec/G395 and MIRI LRS instruments, provided their mixing ratios exceed 10-6 (Gasman et al., 2022). This suggests that the detection of hydrocarbons in the atmosphere of K2-18 b is plausible if methane is abundant and predominant over carbon dioxide, as predicted by Wogan et al. (2024); Schmidt et al. (2025); Jaziri et al. (2025); Hu et al. (2025), and if photochemistry is sufficiently active. The 3.4 and 6.9 µm features observed in the NIRSpec and MIRI data could plausibly be attributed to C2H6 (Luque et al., 2025), providing an alternative to their attribution to DMS (dimethyl sulfide) or DMDS (dimethyl disulfide) by Madhusudhan et al. (2025). In summary, two-atom hydrocarbons are good candidates for detection with current JWST observations and future high-resolution ground-based observations, both in terms of expected abundance and photochemical plausibility. It is necessary to consider them before considering more complex species with a variety of potential sources.
4.3.2 Oxidized organics
Our complementary experimental approach combining MS and IR spectroscopy enables robust identification of H2CO, CH3OH, and CH3CHO formation in oxidized gas mixtures. Jaziri et al. (2025) predict that H2CO, CH3OH, and CH3CHO could reach volume mixing ratios of up to a few parts per million (10-6) in the upper atmosphere of K2-18 b. However, detectability studies with JWST indicate that species such as H2CO and CH3OH require volume mixing ratios exceeding 10-5 to produce measurable features in H2-dominated atmospheres (Huang et al., 2022; Zhan et al., 2022). Their detection in the atmosphere of K2-18 b therefore seems unlikely given current sensitivity limits. Nevertheless, our results suggest that in other H2-rich exoplanets with a higher CO and/or CO2 content, the formation of oxidized organic species may be particularly efficient. In such environments, these molecules could form more abundantly and potentially reach detectable levels with JWST or with future high-resolution ground-based instruments on the Extremely Large Telescope (ELT), such as ANDES (ArmazoNes high Dispersion Echelle Spectrograph, Palle et al. 2025) and PCS (Planetary Camera and Spectrograph, Kasper et al. 2021).
5 Conclusion
We studied out-of-equilibrium chemistry in H2-dominant atmospheres typical of temperate sub-Neptunes by combining experimental and numerical simulations. This approach complements previous laboratory studies by enabling more accurate molecular identification and clearer elucidation of reaction pathways. Our experiments, which are based on the combined use of MS and IR spectroscopy, enabled the reliable detection of reduced (C2H2, C2H4, C2H6) and oxidized (CO, CO2, H2O, H2CO, CH3OH, CH3CHO) species in gas mixtures covering a wide range of C/O ratios and metallicities. The use of the 0D photochemical model allowed us to extract abundance trends and identify formation pathways.
Hydrocarbon compounds are produced in methane-rich atmospheres in particular. The efficiency of this organic growth increases with metallicity in CH4-rich and CO-rich atmospheric analogs. In CO2-rich atmospheric analogs, oxidative loss reactions become predominant and limit the formation of hydrocarbons. The formation of hydrocarbons from methane can be explained by reaction pathways that have already been well studied in the atmospheres of the Solar System, whereas their formation from oxidized reactants — CO or CO2 — involves more reaction intermediates and a change in the predominance of certain pathways. This explains the lower abundance of reduced organic compounds in an oxidized atmosphere, combined with the emergence of oxidation pathways that consume reaction intermediates, which have more influence for low C/O ratios. The addition of oxygen diversifies the chemistry, and in CO-rich and CO2-rich mixtures in particular, it leads to the formation of the oxidized organic compounds presented above. The combination of CH4 and CO has a twofold effect. On the one hand, it provides carbon-rich atmospheres, which promote the efficient formation of hydrocarbons. On the other hand, it leads to moderately oxygenated atmospheres, which promote the formation of oxidized organic compounds without overly enhancing the pathways for hydrocarbon destruction.
Our study highlights the importance of adopting a synergistic approach combining experiments, modeling, and observations in order to deepen our understanding of exoplanet atmospheres. Each field offers unique perspectives, but through close collaboration their potential is greatly enhanced, and therefore strengthening these interdisciplinary efforts is essential.
6 Perspectives
The gas mixtures studied here cover a wide range of potential atmospheres. However, these are still fairly simple mixtures at a fixed temperature, which do not cover all possible atmospheric compositions of exoplanets. For example, the effect of temperature, other oxygen-containing reactants such as water in higher proportions, nitrogenous species, sulfur species, could be studied in the future. It would be very useful to study the detectability of photoproducts in detail, both for JWST observations and for future high-resolution ground-based observations.
Data availability. Mass spectra, IR spectra, and 0D model data can be downloaded from Zenodo: https://doi.org/10.5281/zenodo.17868179.
Acknowledgements.
This project has received funding from the European Research Council (ERC) under the ERC OxyPlanets projects (grant agreement No. 101053033). We would like to thank the anonymous referee for their excellent comments, which improved the presentation of our results. The authors thank Olivier Guaitella for kindly lending the multipass infrared cell, which was essential for this work. We would like to extend our warmest thanks to Joleen Csuka and Elisha Wesley Belibio for their valuable and helpful advice on proofreading and copyediting.
Appendix A Experimental identification of species
We report here the MS fragmentation patterns of hydrocarbon compounds with two and three atoms (Figure 11), the reference IR spectra from the HITRAN database (Figure 12) and the main MS and IR features used to identity products (Table 2).
Figure 11: Mass spectrum of an irradiated gas mixture of 95% H2 + 5% CH4 (Gray). Fragmentation patterns (FP) of the two-carbon atoms (11(a)) and three-carbon atoms (11(b)) compounds (Pink for the alkanes, orange for the alkenes, green for the alkynes). Each FP is normalized to the intensity of the experimental spectrum at the m/z value corresponding to the most intense peak of the FP. 95% uncertainty bars are displayed in black.
Figure 12: Infrared features calculated from the HITRAN database for molecules with up to 2 carbon atoms studied in this study, with a fixed composition of 1% for each of them in a gas mixture of 1 mbar at 340 K.
Table 2: Features of products identified by MS and IR spectroscopy.
Molecules Main MS features (u) Main IR features (cm-1) Main IR features (µm)
C2H2 26 - 25 - 24 - 27 - 28 730 [p] 13.70 [p]
C2H4 28 - 27 - 26 - 25 - 24 950 [p] 10.53 [p]
C2H6 28 - 27 - 30 - 29 - 26 2900-¿3000 [s] - 1300-¿1600 [s] 3.45-¿3.33 [s] - 7.69-¿6.25 [s]
CH2C2H2 40 - 39 - 38 - 37 - 36 1910-¿1990 [s] 5.24-¿5.03 [s]
CH3C2H 40 - 39 - 38 - 37 - 36 633 [p] - 2100-¿2200 [s] 15.80 [p] - 4.76-¿4.55 [s]
C3H6 41 - 39 - 42 - 27 - 40 912 [p] 10.96 [p]
C3H8 29 - 28 - 27 - 44 - 43 2900-¿3000 [s] - 1300-¿1600 [s] 3.45-¿3.33 [s] - 7.69-¿6.25 [s]
H2CO 29 - 30 - 28 1745 [p] 5.73 [p]
CH3OH 31 - 32 - 29 - 15 1033 [p] 9.68 [p]
CH3CHO 29 - 44 - 43 - 15 1745 [s] 5.73 [s]
[p] : single peak, narrow       [s] : set of peaks, broad